a rectangular field is to be emclosed by a fence and divided into three lots by fences parallel to one of the sides. find the dimensions of the largest field that can be enclosed with 800 feet of fencing


1
Expert's answer
2021-01-08T11:40:17-0500

Solution:

"2L+4w=800"

so "L=400-2w"


"A=L*w"

"A=(400-2w)w"

"A=-2w^2+400w"

"\dfrac{dA}{dw}=0 \;"

"=-4w+400"

"w=100"

"L=400-200=200"


Answer:

Dimensions of the largest field that can be enclosed with 800 feet of fencing are 200, 100.